### v2

# Social Genome: Grounded Social Reasoning Abilities of Multimodal Models

Abstract
Social reasoning abilities are crucial for AI systems to effectively interpret and respond to multimodal human communication and interaction within social contexts. We introduce Social Genome, the first benchmark for fine-grained, grounded social reasoning abilities of multimodal models. Social Genome contains 272 videos of interactions and 1,486 human-annotated reasoning traces related to inferences about these interactions. These traces contain 5,777 reasoning steps that reference evidence from visual cues, verbal cues, vocal cues, and external knowledge (contextual knowledge external to videos). Social Genome is also the first modeling challenge to study external knowledge in social reasoning. Social Genome computes metrics to holistically evaluate semantic and structural qualities of model-generated social reasoning traces. We demonstrate the utility of Social Genome through experiments with state-of-the-art models, identifying performance gaps and opportunities for future research to improve the grounded social reasoning abilities of multimodal models.
Social Genome:
Grounded Social Reasoning Abilities of Multimodal Models
Leena Mathur*†1, Marian Qian*1, Paul Pu Liang2, Louis-Philippe Morency1 Carnegie Mellon University1, Massachusetts Institute of Technology2 {lmathur,marianq,morency}@cs.cmu.edu, ppliang@mit.edu
11footnotetext: equal contribution, †corresponding author
1 Introduction
Humans rely on social reasoning to interpret and navigate everyday interactions Gagnon-St-Pierre et al. (2021). This form of reasoning is a core competency of social intelligence Kihlstrom and Cantor (2000); Conzelmann et al. (2013), occurs with specialized neural and cognitive systems Cao et al. (2024); Read et al. (2013), and involves integrating information over time from multimodal behaviors such as gestures, language, and prosody Morency (2010); Read and Miller (2014); Liang et al. (2024). Multimodal cues are often fine-grained (e.g., a fleeting glance), interleaved (e.g., a shrug followed by a sigh), and context-dependent, requiring external knowledge of contextual information to be interpreted accurately Hechter and Opp (2001).
Figure 1: Reasoning over multimodal social interactions involves extracting, integrating, and referencing evidence from multiple behavioral modalities, as well as information from external knowledge.
Developing algorithms for multimodal social reasoning will be essential to advance artificial intelligence (AI) systems with social intelligence Mathur et al. (2024). When AI systems reason about human social interactions, it is important for systems to have the ability to generate interpretable explanations with accurate, grounded references to fine-grained multimodal behaviors and external social knowledge concepts that inform inferences. Figure 1 visualizes these aspects of multimodal social reasoning. This capability is especially important for AI systems reasoning about human social interactions in high-stakes domains, such as healthcare agents and assistive robots.
Figure 2: A sample reasoning trace from the Social Genome benchmark. Reasoning traces in Social Genome contain fine-grained, multimodal social cues and references to external knowledge informing the social inference. Social reasoning traces produced by humans can contain complex reasoning paths (sample visualized above) that reference and build upon multimodal evidence and external knowledge across temporal segments of interactions.
Progress towards improving multimodal social reasoning abilities of models has been limited by a lack of evaluation tasks – measuring a capability is an essential first step towards advancing it. We introduce Social Genome, the first benchmark for grounded multimodal social reasoning that includes 272 videos of face-to-face interactions and 1,486 human-annotated reasoning traces explaining inferences about social information in these videos. Across these traces, Social Genome contains 5,777 social reasoning steps. Each reasoning step is tagged with the modality of information being referenced: visual, verbal, and vocal cues from social interactions in videos, and external knowledge of contextual information that human annotators used to perform social inferences (information external to stimuli in videos). Reasoning traces in Social Genome reference over 11,000 entities (people, objects, concepts), over 5000 multimodal cues, and over 2,900 external knowledge observations. Social Genome is the first social reasoning benchmark to include external knowledge and dense reasoning traces. A sample human-annotated reasoning trace is provided in Figure 2.
This paper defines metrics to holistically assess semantic and structural aspects of model-generated social reasoning traces. We demonstrate the utility of Social Genome by using these metrics to distill insights regarding social reasoning capabilities and limitations in state-of-the-art (SOTA) multimodal models. For example, we find that models struggle to perform well under both zero-shot and in-context learning (ICL) settings, demonstrating the significant challenge of building this understudied form of reasoning in models. Our findings contribute novel insights regarding gaps and opportunities for future research to improve the grounded social reasoning abilities of multimodal models.
2 Background
Prior research on social reasoning in models has primarily focused on the ability of models to interpret text-based social scenarios and perform question-answering (QA) tasks about characters’ motivations, intents, and actions; Social IQa remains a key unimodal benchmark in this area Sap et al. (2019). SOTA language models can accurately perform a majority of the inferences in Social IQa, but a gap remains between model and human performance Sap et al. (2022); Shapira et al. (2024). SOTA models have also struggled with text-based QA tasks that probe competencies relevant to social reasoning, specifically theory-of-mind to interpret the goals and beliefs of characters Le et al. (2019); Shapira et al. (2024); Ullman (2023); Kim et al. (2023). Crowd-sourced knowledge bases of norms Forbes et al. (2020); Ziems et al. (2023) have been useful to inform social reasoning research.
The ability of models to perform social reasoning on multimodal human interactions, in particular face–to–face, embodied, real-world social interactions, has been comparatively understudied. Key benchmarks include the video QA tasks of Social-IQ 1.0 Zadeh et al. (2019) and Social-IQ 2.0 Wilf et al. (2023); both examine model QA accuracy when answering questions about social interactions in videos. SOTA models have struggled to perform well on Social-IQ 2.0 Xie and Park (2023); Pirhadi et al. (2023); Li et al. (2024c); Agrawal et al. (2024); Chen et al. (2024). The community’s prior focus on QA accuracy to assess social reasoning ability has not enabled researchers to study the extent to which models can effectively reference multimodal cues and external knowledge informing inferences. Models with high accuracy on QA tasks can perform poorly at generating valid or comprehensive reasoning traces Jhamtani and Clark (2020); Gu et al. (2023). Model reasoning trace quality has not been studied before in the context of multimodal social reasoning. We introduce Social Genome as the first benchmark to study grounded, fine-grained social reasoning abilities of multimodal models.
3 Building Social Genome
3.1 Sourcing Seed Videos and Questions
The Social Genome benchmark contains 272 seed videos and 1486 questions adapted from the publicly-available Social-IQ 2.0 dataset Wilf et al. (2023) (details on data sourcing in Appendix A.1). Videos include real-world face-to-face dyadic and multi-party interactions (1 minute per video, $\sim$4.5 total hours), and questions probe social dynamics, behaviors, emotions, and cognitive states of both individuals and groups. From seed videos and questions, Social Genome introduces a new set of 1486 human reasoning traces with 5700+ steps that answer these questions.
3.2 Task Notation
Given a video $V$ depicting a social interaction, a question $Q$ about social content in the interaction, and a corresponding set of answer options $A=\{A_{\text{correct}},A_{\text{incorrect}_{1}},A_{\text{incorrect}_{2}},A_{\text{incorrect}_{3}}\}$, a model performing the Social Genome task must generate a reasoning trace $R=\{e_{1},e_{2},\dots,e_{n}\}$, where each reasoning step $e_{i}$ represents a single piece of evidence contributing toward the social inference to select an answer $A_{\text{a}}$ from $A$. Each reasoning step $e_{i}$ must be tagged with two attributes: (1) a modality tag $m_{i}\in\{\textit{visual, verbal, vocal, }n/a\}$ indicating the communication modality of the evidence and (2) an external knowledge tag $k_{i}\in\{\text{{yes}, {no}}\}$, indicating whether the evidence references external knowledge of contextual information. This task to generate $\mathit{R}$ and answer question Q evaluates a model’s ability to extract and reference multimodal aspects of human communication and knowledge informing social inferences. Given the input tuple $(V,Q,A)$, each model performing the Social Genome task will produce an output tuple $(A_{\text{a}},R)$. Metrics in Social Genome study the social inference accuracy of $A_{a}$ and the semantic and structural aspects of social reasoning in $R$.
3.3 Social Reasoning Trace Annotations
Human Annotation
Given a video $V$, question $Q$, and answer options $A$, annotators read $Q$ and $A$, watched $V$, and wrote reasoning trace $R$. Annotations were collected with an IRB-approved Prolific study (details in Appendix A.2).
Grounded and Fine-Grained Behaviors
Humans perceive and build upon low-level observations of fine-grained behaviors (e.g, shifts in body language) and high-level, top-down processing (e.g., implicit situational knowledge) when interpreting social scenes Baird and Baldwin (2001); Bodenhausen and Morales (2013). Annotators were instructed to reference any low-level and high-level evidence they relied upon to answer questions: for example, low-level evidence might be "the woman takes a step back with her mouth wide open (visual cue)" and high-level evidence that interprets that low-level cue might be "the woman is surprised (external knowledge regarding how ’surprise’ might manifest"). For each step, annotators tagged the perceptual modality referenced (visual, verbal, vocal, n/a).
Grounded External Knowledge
Annotators tagged each reasoning step with $yes$ or $no$ to indicate whether external knowledge was referenced. External knowledge includes contextual norms, cultural expectations, and prior understanding of social commonsense Forguson and Gopnik (1988) that goes beyond stimuli in the video. For example, if a man raises his arm and the annotator recognizes his movement as a "high five," the identification of that gesture and its interpretation as "triumphant" is based on external knowledge of social norms.
Ensuring Annotation Quality
Trained experts validated each Prolific annotation. They watched each video, read each QA tuple, and read the annotation to ensure that traces represented valid reasoning, had correct modality and external knowledge tags, and referred to relevant information. Cases of incomplete annotation or deviation from instructions were fixed, as discussed in Appendix A.3.
3.4 Dataset Statistics
Social Genome contains 1486 detailed human-annotated social reasoning traces with 5,777 total steps, 3.89 $\pm$ 1.68 steps per trace (minimum of 1 step and maximum of 10 steps), 43 $\pm$ 26 words per trace, and 11$\pm$ 5 words per step. Reasoning steps draw on multimodal evidence, with 44% of steps referencing visual cues, 27% referencing verbal cues, and 17% referencing vocal cues. Overall, 77% of traces referenced at least one visual cue, 63% referenced at least one verbal cue, and 47% referenced at least one vocal cue.
External knowledge plays a critical role in Social Genome: 51% of reasoning steps referenced external knowledge, with each trace referencing an average of 2 pieces of external knowledge. With spaCy named entity recognition (NER) Honnibal et al. (2020), we found 11,253 entities (people, objects, concepts) mentioned, with 7.6 unique entities and 2.23 emotions referenced per reasoning trace, demonstrating the high density of annotations.
3.5 Social Reasoning Metrics and Statistics
We develop a set of metrics to holistically evaluate semantic and structural aspects of social reasoning traces generated by models performing tasks in the Social Genome benchmark. Collectively, these metrics reveal strengths and weaknesses in model social reasoning and multimodal grounding abilities and the extent to which model traces differ from human reasoning. For each sample, we compute the following metrics between model reasoning trace $\mathit{R_{M}}$ with $n$ steps $e_{1}...e_{n}$ and the corresponding human trace $\mathit{R_{H}}$ with $m$ steps $h_{1}...h_{m}$.
Accuracy
Measures accuracy of the model-generated answer by comparing it to ground truth. Human annotator accuracy on Social Genome is 0.85 (Appendix A.4). Higher values indicate stronger social inference ability (max value is 1).
Similarity-Trace ($S_{\text{trace}}$)
Measures the high-level semantic similarity between $\mathit{R_{M}}$ and $\mathit{R_{H}}$:
$S_{\text{trace}}=\frac{\langle\mathbf{R_{M}},\mathbf{R_{H}}\rangle}{\|\mathbf{R_{M}}\|\cdot\|\mathbf{R_{H}}\|}$
where $\mathbf{R_{M}}$ is the aggregate embedding of evidence steps in $\mathit{R_{M}}$ and $\mathbf{R_{H}}$ is the aggregate embedding of evidence steps in $\mathit{R_{H}}$. The embedding model all-MiniLM-L6-v2 Reimers and Gurevych (2019), selected for its efficiency and accuracy, was used to embed evidence steps for this and other semantic similarity metrics. Higher values indicate stronger alignment in semantic information between $R_{M}$ and $R_{H}$ (max value is 1).
Similarity-Step ($S_{\text{step}}$)
Measures the fine-grained semantic similarity between $\mathit{R_{M}}$ and $\mathit{R_{H}}$. For each step $e_{i}$ in $\mathit{R_{M}}$, the metric identifies its closest semantic step $h_{j}$ in $\mathit{R_{H}}$. The final metric is the mean of these maximum similarity values:
$S_{\text{step}}=\frac{1}{n}\sum_{i=1}^{n}\max_{j}\frac{\langle\mathbf{e_{i}},\mathbf{h_{j}}\rangle}{\|\mathbf{e_{i}}\|\cdot\|\mathbf{h_{j}}\|}$
where $\mathbf{e_{i}}$ is the embedding of evidence step $e_{i}$ and $\mathbf{h_{j}}$ is the embedding of evidence step $h_{j}$. Higher values reflect stronger alignment in semantic information between fine-grained steps of evidence in $R_{M}$ and $R_{H}$ (max value is 1).
Similarity-Num Steps ($S_{\text{num}}$)
Measures the number of reasoning steps in $\mathit{R_{M}}$ with a semantic similarity above a threshold $\tau$, when compared to any step in $\mathit{R_{H}}$. This metric is computed as:
$S_{\text{num}}=\sum_{i=1}^{n}\mathbbm{1}\left(\max_{j}\frac{\langle\mathbf{e_{i}},\mathbf{h_{j}}\rangle}{\|\mathbf{e_{i}}\|\cdot\|\mathbf{h_{j}}\|}>\tau\right)$
where $\mathbbm{1}(\cdot)$ is the indicator function and $\tau=0.6$ (empirically selected). Higher values indicate more semantically-aligned evidence between $R_{M}$ and $R_{H}$ (max value is $n$).
DifferenceSequence ($DS$)
Measures structural similarity between $\mathit{R_{M}}$ and $\mathit{R_{H}}$ using the respective modality sequences $\mathit{S_{M}}$ and $\mathit{S_{H}}$ from the reasoning traces (e.g., ["visual", "external knowledge"]). A similarity score based on edit distance, adapted from the Levenshtein distance, is computed between $\mathit{S_{M}}$ and $\mathit{S_{H}}$ (Appendix B.2). Higher values of $DS$ indicate greater structural similarity between $R_{M}$ and $R_{H}$ (max value is 1).
EmotionMetric
Measures the alignment of emotional content in $\mathit{R_{M}}$ and $\mathit{R_{H}}$. This metric extracts sets of emotions referenced by $\mathit{R_{M}}$ and $\mathit{R_{H}}$ (instructions in Appendix B.3) and computes the overlap in sets. Higher values can indicate stronger emotional alignment between $\mathit{R_{M}}$ and $\mathit{R_{H}}$ (max value is emotion set size in $\mathit{R_{H}}$).
All Modality Steps
Measures the overlapping number of unique modalities in $\mathit{R_{M}}$ and $\mathit{R_{H}}$. Higher values indicate that $\mathit{R_{M}}$ had more overlap with modalities referenced by $\mathit{R_{H}}$ (max value is the number of unique modalities in $\mathit{R_{H}}$).
Visual Steps
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with visual evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to the presence of visual evidence (max value is number of visual steps in $\mathit{R_{H}}$).
Verbal Steps
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with verbal evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to the presence of verbal evidence (max value is the number of verbal steps in $\mathit{R_{H}}$).
Vocal Steps
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with vocal evidence to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to the presence of vocal evidence (max value is the number of vocal steps in $\mathit{R_{H}}$).
External Knowledge Steps
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with external knowledge evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to external knowledge references (max value is the number of external knowledge steps in $\mathit{R_{H}}$).
NumSteps ($NS$)
Measures the absolute difference in the number of reasoning steps between $\mathit{R_{M}}$ and $\mathit{R_{H}}$. Lower values indicate stronger alignment in length between model and human chains (value of 0 indicates that $\mathit{R_{H}}$ and $\mathit{R_{H}}$ are the same length).
Figure 3: Performance of models across number of few-shot ICL samples ($k$) and ground truth noted in gray.
With these metrics, Social Genome enables a holistic study of multimodal, grounded social reasoning.
The first six metrics focus on social inference accuracy, semantic similarity, and structural similarity between model and human reasoning traces. The final six metrics focus on fine-grained multimodal evidence and external knowledge referenced by models, in comparison to evidence referenced by humans.
3.6 Social Genome ICL Training Set
To create samples for ICL experiments, we randomly sampled 16 questions from unique videos in the training set of Social-IQ 2.0 and collected reasoning trace annotations in the same format as annotations in Section 3.3. ICL experiments in Section 4 are conducted by providing each model with different numbers of training samples $k\in\{0,1,2,4,8,16\}$, before the model is given an input tuple ($V$, $Q$, $A$) and generates a sequence of tokens with an answer $A_{a}$ and reasoning trace $R$.
4 Social Reasoning Experiments
We use Social Genome to study the performance of multimodal video understanding models in fine-grained, grounded social reasoning. These models exhibited SOTA performance on video understanding tasks and take videos as input. We tested 2 closed-source models and 5 open-source models: Gemini-1.5-Flash Team et al. (2024), GPT-4o OpenAI , LLaVA-Video and LLaVA-Video-Only Zhang et al. (2024c), LongVA Zhang et al. (2024a), Video-ChatGPT Maaz et al. (2023), and VideoChat2 Li et al. (2023a). Models have different architectures, pretraining data, and fine-tuning tasks, and models generated reasoning traces and answers for all samples (details in Appendix C). LLaVA-Video-Only answered questions, but did not generate reasoning traces that could be studied; this model does not appear in trace-related metrics.
4.1 Quantitative Results and Insights
Figure 3 visualizes each model’s average performance for our 12 metrics [^1] Visualized metrics are normalized to [0,1] by human baselines, except for those with an upper bound of 1 by definition (Accuracy, Similarity-Trace, Similarity-Step, DifferenceSequence) and absolute measures (NumSteps)., with human baselines in gray. Results tables (Tables 3, 4, 5) are in Appendix C. Key findings are discussed below.
Social Inference Accuracy
Human social inference ability is substantially higher than all models, as seen in results from the Accuracy metric. Gemini-1.5-Flash and GPT-4o achieve the highest accuracies of 74.4% and 71.0% respectively ($k$ = 0), approximately 10-15% lower than human annotator accuracy (85.3%), answering $\sim$30of questions incorrectly. Closed-source models outperform open-source models in social inference. The highest-performing open-source model was LLaVA-Video at 62.9% ($k$ = 0). Gemini-1.5-Flash and GPT-4o are much larger than open-source models, suggesting that increased model scale is useful, but not sufficient, for accurate social inference. Video-ChatGPT and VideoChat2 perform 30-40% lower than other open-source models across all values of $k$. These models have the smallest context length and process the fewest frames, constraints which may influence performance.
Social inference accuracy for models decreased as the number of few-shot ICL samples increased, with the exception of GPT-4o, which demonstrated a slight improvement at $k$ = 16. Few-shot ICL conditions language models on tasks by providing examples of inputs and outputs Liu et al. (2024) and can be viewed as a form of inductive reasoning, as the model is tasked with inferring generalizable rules from a set of examples. This technique has improved model reasoning abilities in domains such as mathematics and code generation Dong et al. (2024); Zhou et al. (2022); Patel et al. (2023), which have explicit rules and formal structure Galotti (1989). In contrast to these domains, social reasoning often operates with implicit rules, less formal structure Perkins (1989) and ambiguity in premises Mathur et al. (2024). Our findings suggest that few-shot ICL may not be an effective approach to elicit multimodal social reasoning abilities. Additional experiments (provided in Appendix D) demonstrate that chain-of-thought prompting did not improve model accuracy, and models struggled to perform inferences relying on implicit and contextual knowledge. Our finding from Social Genome that few-shot ICL is insufficient to elicit multimodal social reasoning aligns with findings from unimodal experiments on Social-IQa and ToMi datasets Le et al. (2019); Sap et al. (2019, 2022); Kim et al. (2023).
Semantic Alignment
It is challenging for models to generate social reasoning traces with high semantic alignment to human reasoning, as seen in the results from the Similarity-Trace, Similarity-Step, and Similarity-Num Steps metrics. For Similarity-Trace, only Gemini-1.5-Flash achieved slightly above 50% ($k$ = 0), and LongVA outperformed Gemini-1.5-Flash after $k$ = 4. For Similarity-Step, no models achieved above 50%, but Video-ChatGPT outperformed all models for $k\in\{0,1,2,4\}$ (5-12% higher than GPT-4o and 2-5% higher than Gemini-1.5-Flash). For Similarity-Num Steps, Gemini-1.5-Flash achieved the highest performance (27%), far below ground truth. Low performance can be explained by failure to reference cues that humans reference (“Fine-Grained Grounding" metrics below). Performance on these metrics did not improve as $k$ increased.
Structural Alignment
Model reasoning traces tend to reference multimodal evidence in different amounts and orders than humans, as seen in the results from the DifferenceSequence metric. Model-generated sequences of multimodal evidence that were most structurally-aligned with human sequences were from Gemini-1.5-Flash (0.53 at $k$ = 4) and LLaVA-Video (0.49 at $k$ = 4), both far below the maximum alignment value (1). As $k$ increased, DifferenceSequence metric scores increased for Gemini-1.5-Flash, LLaVA-Video (up to $k$ = 4), and GPT-4o (after $k$ = 2). These findings suggest that structured samples of human social reasoning, as introduced by Social Genome, can be useful when conditioning models to generate reasoning traces with more human-like structure.
Emotion Alignment
It is challenging for models to generate social reasoning chains with high emotional alignment with human reasoning, as seen in the results from the EmotionMetric. All models achieved less than 22%, far below ground truth. While the highest scores were achieved by Gemini-1.5-Flash and GPT-4o at $k$ = 0, scores from several models steadily improved after $k$ = 2.
Figure 4: Representative social reasoning traces from a Human Annotator, Gemini-1.5-Flash, and LLaVA-Video. These examples illustrate the grounded social reasoning abilities and reasoning structures across humans and models.
Fine-Grained Multimodal Grounding
The final 6 metrics study evidence referenced by model reasoning traces. The AllModality Steps metric serves as a proxy for how well models refer to fine-grained, multimodal cues and external knowledge. As seen in AllModality Steps results, all models referenced fewer pieces of multimodal evidence and external knowledge than human reasoning.
While we discuss modality-specific findings below, our use of Social Genome to study SOTA models is currently limited by their varying abilities. The Gemini-1.5-Flash API reports that audio is processed in-parallel on the backend, but other models do not process audio. However, despite the absence of audio, models referenced verbal and vocal evidence inferred from visual frames. For example, LongVA generated vocal evidence about a woman’s "tone suggesting frustration," after generating visual evidence about the woman’s face appearing dissatisfied. As new models are developed in the coming years with abilities to jointly process video and audio, Social Genome will continue to be applicable to study model abilities in grounded multimodal social reasoning.
Visual Steps: Closed-source models exhibited a strong ability to reference visual evidence, with GPT-4o and Gemini-1.5-Flash closer to ground truth than open-source models. Performance on this metric for all models was highest at $k$ = 0.
Verbal Steps: The ability of models to reference verbal evidence showed substantial variation. Gemini-1.5-Flash exhibited the strongest ability to reference verbal cues (0.90 at $k=0$). Model performance on this metric did not improve as $k$ increased, and GPT-4o referenced substantially fewer verbal cues in comparison to other models.
Vocal Steps: The ability of models to reference vocal evidence was substantially lower than human ability. LongVA referenced more vocal evidence than other models at $k$ = 0. Model performance on this metric did not improve as $k$ increased.
External Knowledge Steps: The ability of models to reference external knowledge in social reasoning traces was lower than humans. In contrast to trends observed in other modality step metrics, we find that providing additional few-shot samples improved the ability of Gemini-1.5-Flash, GPT-4o, LongVA (up to $k$ = 4), and LLaVA-Video (up to $k$ = 4) to reference external knowledge. These findings suggest that human social reasoning traces can be used to condition models to ground social reasoning with external knowledge references.
Num Steps: Model reasoning traces varied in length and contained more steps than human traces. LLaVA-Video and LongVA generated model reasoning traces that were most aligned with the length of human traces. Providing additional few-shot samples improved the ability of models to align with human social reasoning trace length.
4.2 Human and Qualitative Evaluation
We conducted human evaluation of model reasoning traces. Trained annotators analyzed 48 samples from all models, with model names and $k$ values anonymized. Traces were rated to assess references to low-level cues (fine-grained), information cross-referenced across steps (compositional), relevant evidence (comprehensive), correctness of modality tags, and validity of reasoning (criteria are in Appendix E). Annotator agreement was computed with Cohen’s Kappa $\kappa$ Cohen (1960).
Reasoning traces from Gemini-1.5-Flash and GPT-4o received the highest ratings for fine-grained ($\kappa$ = 0.87), compositional ($\kappa$ = 0.92), comprehensive ($\kappa$ = 0.94), and valid ($\kappa$ = 0.94) reasoning, followed by LLaVA-Video and LongVA, with VideoChatGPT and VideoChat2 rated lowest. These findings validate model performance trends observed in Section 4.1. Modality tag correctness across samples was 98%, supporting the validity of the Social Genome automated modality metrics and reasoning trace processing (Appendix B.4).
We visualize representative traces in Figure 4 from a human annotator, Gemini-1.5-Flash, and LLaVA-Video (closed-source and open-source models with highest inference accuracy at $k$ = 0).
Evidence and Error Propagation
This figure illustrates the strong ability of Gemini-1.5-Flash to reference and integrate multimodal cues and external knowledge during social reasoning. In contrast, LLaVA-Video did not reference low-level cues and based its reasoning upon an incorrect premise, and this error propagated to an incorrect inference. In addition, the human trace referred to several fine-grained, subtle behaviors (e.g., lip movements) that are not present in the model traces, yet do influence the scene interpretation. Similarity metrics and modality-specific metrics in Section 3.5 serve as automated proxies to estimate these types of information gaps between human and model reasoning, as shown in these qualitative examples. Our findings motivate future work on model training data, training approaches, and architectures that better capture fine-grained cues, as well as algorithms to handle error propagation in social reasoning.
Hierarchical Social Reasoning
Figure 4 qualitatively illustrates the hierarchical structure of human social reasoning traces, in which low-level cues (e.g., brief lip movement) are referenced, combined, and re-interpreted as intermediate evidence for further reasoning. This “forking" reasoning structure is common for humans interpreting everyday situations, unlike the linear “long chain" structure for formal reasoning in domains such as mathematics Perkins (1989); Galotti (1989). As seen in Figure 4, compared to human traces, model traces have flatter structures with less hierarchical composition of social cues. Flatter structures have the potential to treat cues as isolated evidence and overlook intermediate evidence in reasoning paths. Our findings motivate future work to train models capable of more hierarchical social reasoning.
5 Conclusion
This paper introduces Social Genome, the first benchmark for fine-grained, grounded social reasoning abilities of multimodal models. The reasoning traces contributed by Social Genome include references to multimodal cues and external knowledge concepts that humans find useful when performing social inferences. We define metrics to assess semantic and structural aspects of reasoning traces, and we contribute novel insights regarding gaps and opportunities to improve the grounded social reasoning capabilities of multimodal models.
Future AI systems interacting with humans in face-to-face, embodied contexts must have the ability to reason about social interactions and ground reasoning in concrete multimodal evidence and external knowledge concepts. Social Genome serves as a first step towards studying and advancing this form of reasoning in AI systems.
6 Limitations
Social Reasoning in Natural Language
The current scope of Social Genome focuses on studying model-generated social reasoning traces in natural language. This scope is necessary and relevant to contexts that require AI systems to generate natural language explanations of social inferences – for example, a healthcare agent or hospital robot reasoning about human nonverbal behaviors during a nurse-patient social interaction. However, an open question remains regarding the extent to which natural language can effectively represent the nuances of human social interactions and social reasoning Mathur et al. (2024). It is possible that both humans and models verbalizing social reasoning through natural language are not fully capturing why they came to certain inferences Turpin et al. (2023). Several lines of work in reasoning have operated in the latent space of models instead of natural language Hao et al. (2024); Geiping et al. (2025), and we believe Social Genome informs and motivates future work to develop techniques to study social reasoning in the latent space.
Video Contexts
The videos in Social Genome each have a length of 1 minute. We believe that findings from Social Genome regarding model social reasoning abilities about interactions in these 1-minute clips motivate future work to curate longer-form video benchmarks. Social interactions during hour-long videos Zou et al. (2025), for example, would contain more complex social dynamics and longer-form temporal dependencies across humans. In addition, all videos in Social Genome have social interactions in English, and all annotators were required to be proficient in English. As social norms can differ across cultures and languages Hershcovich et al. (2022), we believe future research in multimodal social reasoning would benefit from a larger-scale curation of multimodal interaction data across cultural and linguistic contexts.
7 Ethics
Ethical Annotation
We curated annotations from videos in existing publicly available datasets. We hired workers from Prolific to annotate reasoning traces. All workers received fair compensation for their annotation (\$12 per hour, pro-rated). Worker privacy and confidentiality were respected, with no identifiable information stored. Further details on Prolific annotation are in Appendix A.2.
Risks for Social Reasoning in AI
Social reasoning abilities are essential for future AI systems to effectively work with and alongside humans in real-world settings. Social Genome has the potential to support the research community in studying and advancing these capabilities in AI systems. We envision AI systems using social reasoning to enhance human autonomy, health, and well-being – for example, a healthcare agent reasoning about a patient’s nonverbal behaviors, in order to help doctors better support them. However, these technologies exist within a broader societal context, with multiple stakeholders and potential risks in areas such as privacy, surveillance, and manipulation. We support broader academic research and policy efforts to mitigate against misuse and potential harms of socially-intelligent AI.
Acknowledgements
Leena Mathur is supported by the NSF Graduate Research Fellowship Program under Grant No. DGE2140739. This material is based upon work partially supported by National Institutes of Health awards R01MH125740, R01MH132225, and R21MH130767. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors, and no official endorsement should be inferred. Figure 2 includes icon material available from https://icons8.com.
Appendix A Social Genome Dataset Appendix
A.1 Social Genome Data Sourcing
All videos in Social Genome were sourced from publicly-available Social-IQ 2.0 modeling challenge Wilf et al. (2023). We obtained permission from the authors to access the original test set videos and question-answer tuples from this dataset (275 videos and 1514 QA Tuples). Our use of these videos aligns with Social-IQ 2.0 repository’s MIT license and intended research purpose. We chose to use these test set videos as candidate seed videos to build Social Genome because the answers to questions posed about these videos have not been released online, reducing chances of benchmark contamination Xu et al. (2024). At the current stage, we plan to avoid benchmark contamination in Social Genome by not uploading human social reasoning trace annotations online and using them for research purposes.
We note that prior papers that test models on Social-IQ 2.0 have used the validation set, not the test set Xie and Park (2023); Pirhadi et al. (2023); Guo et al. (2023); Li et al. (2024c); Agrawal et al. (2024); Chen et al. (2024). Therefore, we do not directly compare prior works’ validation set performance with our results on the test set.
During manual inspection of each video and QA tuple, we filtered out QA tuples with answer options containing ambiguity – if at least 2 annotators judged a question to have ambiguous answer options (at least two plausibly-correct answers), we discarded the question. The resulting Social Genome set contained 272 videos and 1486 QA tuples. These samples contain face-to-face dyadic and multi-party social interactions and questions that probe understanding of affective states, causal social dynamics, and social events.
A.2 Social Genome Annotation
Annotators were recruited on the Prolific [^2] https://app.prolific.com platform to perform an IRB-approved study with informed consent regarding our intended data use. Screens for annotators were employed to ensure that participants were adults, had access to a desktop to watch the videos and perform the annotation, were fluent in English, were based in the United States, and completed at least a high school education, and had high prior task approval rates on Prolific (97-100% completion). Each annotator watched 2 videos and provided chains for all questions associated with each video (approximately 10 questions total per annotation). Instructions for annotators are listed in Figure 5. Annotators were compensated at \$12 per hour pro-rated. This annotation process was time-intensive, with each human video annotation taking approximately $\sim$25 minutes.
Before running the Prolific study, we ran 4 iterations of annotation instructions with pilot groups to refine instructions. We experimented with instructions that had annotators provide reasoning traces without knowing the correct answer and while knowing the correct answer. We found that providing annotators with the correct answer option did not change the detail, structure, or coherence of the reasoning traces generated. Therefore, we chose to provide annotators with the correct answer option (Figure 5) and focus the large-scale Prolific data collection on obtaining detailed reasoning traces, instead of additional QA answer responses.
Annotation Instructions In this study, you will be watching short videos that contain social interactions. You will be briefly explaining your thought process along several dimensions while answering questions about each video. Your answers will be entered through a Google spreadsheet. Participants in this study must have the ability to watch videos and write logical text responses in English describing their step-by-step reasoning while interpreting videos. If you give your consent to participate in this study, please select "I agree": <I agree> (If the annotator declines, they move to a page that states: “As you do not wish to participate in this study, please return your submission on Prolific by selecting the ’Stop without completing’ button.") In the provided spreadsheet (link below), you will see 2 videos linked in the "Video URL" column and a set of corresponding questions and answers for each video. The yellow-highlighted box indicates the correct answer for each question. Your goal is to do the following: 1. Click the blue URL next to each set of questions in the "Video URL" column. Watch the video and pay attention to the behaviors of people interacting. 2. For each question associated with each video, provide the reasoning chain in order to answer the question: (a) Read the question, read the answer options, and note the correct answer (highlighted in yellow). (b) In the "Evidence" column, describe the reasoning steps you took to reach the correct answer. Put each step in a separate row in the spreadsheet. (If applicable) steps should start with low-level observations (behavioral information from the video OR external knowledge) and build up to higher-level concepts. 3. For each piece of evidence, in the same row, fill out the "Modality" column drop-down menu to indicate whether your evidence is from visual information (facial movements, gestures, anything vision-related, etc), vocal information (e.g., audio/speech), verbal information (e.g., the actual content of words being spoken), or external knowledge (e.g., cultural norms, your own understanding of what a behavior means). 4. Please aim for 5-10 reasoning steps per question (if applicable). We will look at the spreadsheet during our manual review of submissions. All questions must have evidence presented with the answer selected: <link to the spreadsheet>
Figure 5: Sample instructions provided to annotators on Prolific to provide Social Genome reasoning traces.
A.3 Social Genome Validation
After obtaining human-annotated reasoning chains from Prolific, we conduct validation to ensure the validity and correctness of annotations. Two authors manually watched each video, read each question and corresponding set of answer options, and checked whether chains (1) represented valid reasoning paths, (2) had correct evidence tags for visual, verbal, vocal and external knowledge, and (3) comprehensively referred to relevant low-level social information in videos. If an annotator did not complete an annotation in a satisfactory manner (e.g., incomplete reasoning chain, minimal effort), the task was returned to them on Prolific. If annotations could be rapidly fixed (e.g., changing an incorrect modality tag from "visual" to "vocal"), the authors performed this fix themselves without sending the annotation task to the annotator.
A.4 Social Genome Human Accuracy
For each video in Social Genome, two human annotators on Prolific watched the video, read each question, and selected one of four provided answer options. Annotators were paid \$12 per hour (pro-rated) for completing each study. For each question, annotators also had the option to indicate whether or not they were “uncertain" about the answer option they selected, and the authors examined these samples to ensure all QA tuples in Social Genome had correct answers, to avoid the situation in which a sample could have more than one plausibly-correct answer option. Human annotators on Prolific achieved an accuracy of 85.3% on the 1486 questions in Social Genome. Inter-annotator agreement among Prolific annotators was positive (Cohen’s $\kappa$ = 0.60) Cohen (1960), and answer correctness was confirmed independently by authors, as described earlier.
Appendix B Social Genome Metrics Appendix
B.1 Embeddings for Semantic Similarity
The metrics Similarity-Trace, Similarity-Step, and Similarity-Num Steps are computed with embeddings from the all-MiniLM-L6-v2 from Sentence-BERT Reimers and Gurevych (2019). We found this embedding strong, efficient, and useful for our task; as future embedding models are enhanced and released in the coming years, the Social Genome framework allows for the embedding model called by semantic similarity metrics to be updated.
B.2 DifferenceSequence Metric
For model reasoning chain $R_{M}$ with modality sequence $S_{M}$ and human reasoning chain $R_{H}$ with modality sequence $S_{H}$, the DifferenceSequence ($DS$) metric is computed as a normalized similarity score by adapting the Levenshtein distance Levenshtein (1966) between $S_{M}$ and $S_{H}$:
$DS=1-\frac{\text{Levenshtein}(\mathit{S_{M}},\mathit{S_{H}})}{|\mathit{S_{M}}|+|\mathit{S_{H}}|}$
$\text{Levenshtein}(S_{M},S_{H})\text{ is the following:}\hskip 142.26378pt$:
=$=\begin{cases}|S_{M}|,\text{if }|S_{H}|=0\\ |S_{H}|,\text{if }|S_{M}|=0\\ \min\Big{[}\text{Levenshtein}(S_{M}[:-1],S_{H}[:-1])+\delta,\\ \hskip 10.00002pt\text{Levenshtein}(S_{M}[:-1],S_{H})+1,\\ \hskip 10.00002pt\text{Levenshtein}(S_{M},S_{H}[:-1])+1\Big{]}\text{otherwise}\end{cases}$
with $\delta=\mathbbm{1}(S_{M}[-1]\neq S_{H}[-1])$, where $\mathbbm{1}(\cdot)$ returns $1$ if the final elements differ and 00 otherwise. We use an implementation [^3] https://rapidfuzz.github.io/Levenshtein/index.html that treats a substitution as equivalent to one insertion plus one deletion, making the distance effectively an InDel distance. We compute the minimum number of edits that are needed to transform one sequence to another. Higher edit distances indicate that more edits are needed to align sequences (more dissimilarity).
Therefore, the overall $DS$ similarity metric between $S_{M}$ and $S_{H}$ can range from 0 (maximum number of edits required) to 1 (minimum number of edits required). Higher $DS$ values indicate greater structural similarity between the sequences $\mathit{S_{M}}$ and $\mathit{S_{H}}$, with respect to the type and order of modality evidence being referenced.
Model Prompt Template Trace: Having a model provide a reasoning trace in the response (remove the “few-shot examples" for zero-shot experiments at $k$ = 0): Watch the video and answer questions about social information in the video, following the format of the examples below. <FEW-SHOT EXAMPLE with Question and Answer Options> Which of these is the correct answer? Output either A, B, C, or D. Provide the reasoning steps you took to get to this answer. Give 5 to 10 reasoning steps in bullet point form. Tag each bullet point step as visual, vocal, and verbal modality from the frames or external knowledge. Use the format: The correct answer <insert answer>.The correct answer is <FEW-SHOT RESPONSE WITH ANSWER AND REASONING TRACE>. Which of these is the correct answer? Output either A, B, C, or D. <CURRENT QUESTION AND ANSWER OPTIONS> Provide the reasoning steps you took to get to this answer. Give 5 to 10 reasoning steps in bullet point form. Tag each bullet point step as visual, vocal, and verbal modality from the video or external knowledge. Use the format: The correct answer is <insert answer>. No Trace: Having a model answer the question with no reasoning trace (remove the “few-shot examples" for zero-shot experiments at $k$ = 0): Watch the video and answer questions about social information in the video, following the format of the examples below. <FEW-SHOT EXAMPLE with Question and Answer Options> Which of these is the correct answer? Output either A, B, C, or D. Use the format: The correct answer <insert answer>.The correct answer is <FEW-SHOT CORRECT ANSWER, FEW-SHOT REASONING CHAIN>. Which of these is the correct answer? Output either A, B, C, or D. <CURRENT QUESTION AND ANSWER OPTIONS> Use the format: The correct answer is <insert answer>. Note: For GPT-4o prompts, we found that we needed to replace the word video with frames to avoid error messages associated with video processing (e.g., “I am unable to view or analyze video frames directly. However, I can help answer questions based on descriptions or provide general information. Let me know how I can assist you!"). Other models did not have this issue.
Figure 6: Information on model prompts to obtain reasoning traces and inferences from models.
B.3 Emotion Named Entity Recognition
We perform NER with spaCy Honnibal et al. (2020). In spaCy’s NER v3 configuration, we broadly defined an emotion label as a ‘‘description of how a person is feeling". To avoid identifying words like "feels" as entities, we also passed an example into the spaCy NER configuration that explicitly labeled "feels" as not an emotion entity (e.g., ‘‘She feels sad because her friend didn’t come with her"). There are an average of 2.23$\pm$1.63 emotion entities referenced per human reasoning chain.
B.4 Chain Processing for Metrics
Computing metrics requires a standardized format for model-generated reasoning chains. Several model generations (in particular, the generations from open-source models) required processing before metrics could be computed.
We first parsed generations from models by splitting each generation based on its structure, such as the presence of line breaks, numbering, or sentences. For example, if a model generated a multi-sentence response, but did not include line breaks or numbering within the response, we would split this model output by individual sentences (e.g., splitting on the "." character).
We, then, parsed through each step and remove any steps that simply repeated the question or answer choices. We also removed phrases such as "reasoning step", "reasoning", or "the correct answer", as those phrases were often in steps like "The correct answer is A." or "Below are the reasoning steps:". In addition, during in-context learning experiments, we found that some model generations (in particular, GPT) would contain repetitions of sample chains within the model’s response, leading the response to contain 2-3 reasoning chains. We automatically processed these responses by only taking the final reasoning chain out of these multiple chains and checking that this final chain was answering the original question.
Models were tasked with tagging modalities for each step of their generated reasoning chains. To validate this process, we automatically checked whether each step included visual, vocal, verbal or external knowledge tags. However, models sometimes failed to tag modalities for their chains. To automatically handle these cases, we employed GPT-4o-mini Hurst et al. (2024) to tag modalities. We note that the models that needed this additional validation step were VideoChat2, VideoChatGPT, and LLaVA-Video; generations from these models did not have any modality tagging for the majority of their chains. The authors manually inspected a subset of GPT-generated modality tags for model-generated reasoning traces to verify accuracy.
Appendix C Social Genome Model Appendix
C.1 Model Information
Experiments were conducted with multimodal models that were selected for their SOTA performance on various video understanding tasks and have the ability to take a full video as input (2 closed-source models and 5 open-source models): Gemini 1.5 Flash Team et al. (2024), GPT-4o OpenAI , VideoChat2 Li et al. (2023a), Video-ChatGPT Maaz et al. (2023), LLaVA-Video Zhang et al. (2024c), LLaVA-Video-Only Zhang et al. (2024c), LongVA Zhang et al. (2024a). We summarize the models below, and Appendix Table 1 lists characteristics of these models: context length, tokens per frame, training max frame, parameter count, and backbone. Figure 6 describes the prompts given to models. Our benchmark allows experiments with any model that processes multimodal language and video input and outputs text, allowing Social Genome to be used as a benchmark over time to study social reasoning. Experiments were conducted with one A100 GPU.
Model Property VideoChat2 Video-ChatGPT LLaVA-Video LLaVA-Video-Only LongVA GPT-4o Gemini 1.5
Context 2K 2K 32K 32K 224K 128K 1M
Max Frames 16 100 110 110 2000 - -
Parameters 7B 7B 7B 7B 7B - -
Backbone Vicuna-v0 Vicuna-1.1 Qwen2 Qwen2 Qwen2-Extended - -
Table 1: Information about models tested on Social Genome: VideoChat Li et al. (2023a), VideoChat-GPT Maaz et al. (2023), LLaVA-Video Zhang et al. (2024b), LLaVA-Video-Only Zhang et al. (2024b), LongVA Zhang et al. (2024a), GPT-4o OpenAI , Gemini 1.5 Flash Team et al. (2024). All information is completed based on current public reports and repositories.
VideoChat2
The VideoChat2 model Li et al. (2024b) has an architecture with UMT-L vision encoder Li et al. (2023b), QFormer and Vicuna-7B v0 language model, has 7B parameters, can process up to 16 frames, and was trained with instruction tuning on a collection of 34 tasks spanning conversations, captions, visual question-answering, reasoning, and classification.
Video-ChatGPT
The VideoChat-GPT model Maaz et al. (2023) has an architecture built on top of LLaVA, with a CLIP vision encoder Radford et al. (2021) and a Vicuna-7B v1.1 language model. The VideoChat-GPT model can process up to 100 frames, and was trained on video instruction pairs from the VideoInstruct100K dataset.
LLaVA-Video
The LLaVA-Video model LLaVA-Video-7B-Qwen2 Zhang et al. (2024c) has an architecture with a SigLIP SO400M vision transformer and Qwen2 language model, has 7B parameters, can process up to 110 frames, and was trained on mixture of single image, multi-image, and video tasks from the LLaVA-Video-178K and LLaVA-OneVision datasets Li et al. (2024a).
LLaVA-Video-Only
The LLaVA-Video-Only model LLaVA-Video-7B-Qwen2-Video-Only is identical to the LLaVA-Video model, with the exception of the training data Zhang et al. (2024c). LLaVA-Video-Only was solely trained on the LLaVA-Video-178K dataset.
LongVA
The LongVA model LongVA-7B-DPO Zhang et al. (2024a) aligns a unified multimodal transformer (UMT) with QFormer and aligns this visual encoder with a Qwen2 7B language model. LongVA was trained on visual instruction-following datasets and multimodal document data and has a context length of over 200,000 visual tokens; this longer context length was achieved by extending the context length of the language backbone to train on long text samples, before performing multimodal alignment and additional training to transfer this ability to the multimodal domain.
GPT-4o
The GPT-4o model is a closed-source model from OpenAI Hurst et al. (2024). The technical report for GPT-4o refers to this model as “omnimodal" with the ability to accept inputs from text, audio, image, and video and generate outputs with text, audio, and image. The API access to this model supports video frame inputs and text inputs.
Gemini-1.5-Flash
The Gemini-1.5-Flash model is a closed-source model from Google. The API access to this model supports video inputs and text inputs, up to a context length of approximately 1 million tokens. Gemini-1.5-Flash was distilled from the larger Gemini-1.5-Pro sparse mixture-of-experts transformer Team et al. (2024).
C.2 Model Generation Notes
We note observations here on model generations. VideoChat generations for $k\in{0,1,2,4}$ produced full sentences explaining reasoning traces, however the generation quality eroded for $k\in{8,16}$. Several samples from these settings of $k$ were repeated words and short phrases (e.g.,“ the the the...and and and and"). Similarly, VideoChat-GPT generations for $k\in{0,1,2,4}$ produced full sentences explaining reasoning, however the generation quality eroded for $k\in{8,16}$. Samples from these settings of $k$ were repeated short words and letters (e.g.,“or or or, or" and “B ( ( ( ( ( B"). LLaVA-Video generations for $k\in{0,1,2,4,8,16}$ produced full sentences explaining reasoning, however the generations as $k$ increased in $k\in{4,8,16}$ began to answer fewer questions. LLaVA-Video-Only answered questions, but did not generate reasoning traces; this model was discussed in Section 4 solely for the social inference accuracy metric. These model generation challenges were not observed for LongVA or Gemini-1.5. GPT-4o initially generated “I’m sorry, I can’t assist with that." as one of the reasoning trace steps for several questions, before answering the question.
Appendix D Auxiliary Experiments
D.1 Social Inference Accuracy and Reasoning Trace Lengths
We hypothesized that models may perform worse on inferences that primarily rely on implicit cues and contextual knowledge. One proxy for this reliance is the length of a human trace – if humans perform an inference immediately and only need to verbalize one reasoning step, that step was more likely to involve implicit cues with contextual nuances (e.g., rapidly interpreting body language based on external knowledge). For samples with 1 reasoning step, 53% referenced external knowledge in this first piece of evidence, in contrast to 33% of samples with 5 reasoning steps and 20% of samples with 10 reasoning steps.
We examined model social inference performance across samples with different lengths of human reasoning traces, visualized for $k$ = 0 in Figure 7. Overall, multimodal models social inference performance was lower for samples with shorter human reasoning traces and higher for samples with longer human reasoning traces. This trend was observed for both larger closed-source models and smaller open-source models that represent different training data distributions, architectures, and training techniques. For example, Gemini-1.5-Flash and LLaVA-Video achieved accuracies of 70% and 58%, respectively, for samples with the shortest reasoning traces and both achieved 80% for samples with the longest reasoning traces. These results reinforce the perspective that current models (regardless of size) are not sufficient for strong multimodal social reasoning performance in domains requiring more contextual understanding.
Figure 7: Social inference accuracy of models ($k$ = 0) across samples with different numbers of human-annotated reasoning steps (binned into quintile by reasoning trace length).
Figure 8: Social inference of models that generated reasoning traces with answers (Trace) compared to models that generated solely answers (No Trace), across different numbers of few-shot samples $k$.
Table 2: Human evaluation mean annotator scores for reasoning trace samples across models.
Model Shot Fine-Grained Compositional Modality Tags Validity Comprehensive
VideoChat2 0.0 1.50 1.00 0.5 0.0 1.25
1.0 1.00 1.00 1.0 0.0 1.00
4.0 1.00 1.00 1.0 0.0 1.00
16.0 1.00 1.00 1.0 0.0 1.00
Video-ChatGPT 0.0 1.00 1.00 1.0 0.0 1.00
1.0 1.00 1.00 1.0 0.0 1.00
4.0 1.00 1.00 1.0 0.0 1.00
16.0 1.00 1.00 1.0 0.0 1.00
LongVA 0.0 2.75 2.75 1.0 1.0 4.50
1.0 3.00 3.25 1.0 1.0 3.75
4.0 2.25 3.50 1.0 1.0 3.50
16.0 2.25 2.75 1.0 1.0 3.25
LLaVA-Video 0.0 3.00 2.50 1.0 1.0 4.00
1.0 2.00 1.75 1.0 0.5 3.00
4.0 1.00 1.00 1.0 0.0 1.00
16.0 1.00 1.00 1.0 0.0 1.00
GPT-4o 0.0 3.50 3.00 1.0 1.0 5.00
1.0 3.50 3.00 1.0 1.0 5.00
4.0 4.00 3.50 1.0 1.0 4.50
16.0 4.00 4.00 1.0 1.0 4.50
Gemini-1.5-Flash 0.0 4.25 4.25 1.0 1.0 4.75
1.0 4.00 2.50 1.0 1.0 4.75
4.0 3.75 3.75 1.0 1.0 3.00
16.0 4.25 4.25 1.0 1.0 4.50
D.2 Does Chain-of-Thought Prompting Elicit Multimodal Social Reasoning?
Chain-of-thought (CoT) prompting Wei et al. (2022) has been a prevalent approach to elicit model reasoning abilities in domains such as mathematics and code generation Li et al. (2025). We explore the effectiveness of this technique for multimodal social reasoning with Social Genome. Figure 8 visualizes social inference results for models under two settings: Trace (models generate step-by-step CoT traces while answering the question) and No Trace (models generate only answers and do not generate "step by step" traces). We test models with zero-shot and few-shot settings, with $k\in$0, 1, 2, 4, 8, 16}. Prompts are described in Figure 6.
We find that CoT prompting does not substantially improve the social reasoning performance of models, with the exception of GPT-4o (using CoT performs 6-8% higher than without CoT after $k$=4). Unlike domains such as mathematics with more formal step-by-step reasoning paths, social reasoning often involves interpreting and integrating ambiguous, context-dependent cues across actors, modalities, and time Mathur et al. (2024). Social inference is a form of informal reasoning that does not often verbalize as a “chain-like" process Galotti (1989). Models with CoT prompting have been found to rely on task priors from pretraining distributions Chochlakis et al. (2024); it is possible these priors are less effective to elicit social reasoning, compared to other forms of reasoning.
Appendix E Human Evaluation Criteria
Trained annotators analyzed 48 samples from all models at $k\in$ 0, 1, 4, 16. The $k$ values and model identities associated with traces were anonymized before presenting samples to annotators in a spreadsheet to rate. For each reasoning trace, annotators watched the corresponding video, read the question, read the answer options, and read the correct answer, in order to rate the quality of the reasoning trace. Human evaluation was conducted to assess reasoning traces along the following dimensions:
Fine-grained: The extent to which reasoning traces were fine-grained was assessed on a scale of 1 to 5, with 1 for no references to low-level behavior and 5 for dense references (e.g., references to low-level behaviors in a majority of steps).
Compositional: Compositionality in reasoning traces was assessed on a scale of 1 to 5, with 1 for minimal compositionality (no cross-references of information across reasoning steps) and 5 for high compositionality across steps.
Comprehensive: The extent to which reasoning traces were comprehensive was assessed on a scale of 1 to 5, with 1 referring to traces missing critical information (not referencing relevant video content) and 5 being fully-comprehensive (capturing all relevant content towards an inference).
Modality Tag Correctness: The accuracy of modality and external knowledge tags for each of the reasoning steps within a given trace was assessed with a binary score (1 for correct, 0 for the presence of any error in the trace).
Validity of Reasoning: The validity of the reasoning in a trace, referring to whether or not the trace represented logical combinations of information. It is possible for a reasoning trace to reference minimal low-level information, yet still represent valid reasoning (motivating the inclusion of this dimension). This dimension was given a binary rating (1 for valid, 0 for invalid).
Raw averages from this human evaluation process are presented in Table 2. We discuss findings from human evaluation in Section 4.2. Annotator agreement was computed with Cohen’s Kappa $\kappa$ Cohen (1960). We found strong inter-annotator agreement in ratings across dimensions: $\kappa$ = 0.87 for "fine-grained" ratings, $\kappa$ = 0.92 for "compositional" ratings, $\kappa$ = 0.94 for "comprehensive" ratings, and $\kappa$ = 0.94 for "validity of reasoning" ratings. The "modality tag correctness" ratings across samples was 98% with errors specifically occurring in modality tags for VideoChat2 traces.
We note that reasoning trace quality for LLaVA-Video, VideoChat2, and Video-ChatGPT, in particular, decreased at $k$ = 4. These findings from human evaluation are aligned with quantitative findings in Section 4 and subjective model generation observations in Appendix C.2.
Table 3: Social Inference Accuracy on Social-Genome for Models Across Different Numbers of Shots ($k$)
Model Chain Type k = 0 k = 1 k = 2 k = 4 k = 8 k = 16
VideoChat2 Chain 0.2624 0.2779 0.2746 0.2725 0.0000 0.0000
No Chain 0.2712 0.2793 0.2806 0.2611 0.2483 0.0000
Video-ChatGPT Chain 0.3560 0.2645 0.2550 0.2503 0.0000 0.0000
No Chain 0.3762 0.2544 0.2746 0.2806 0.0000 0.0000
LongVA Chain 0.5828 0.4973 0.5283 0.5451 0.5363 0.5565
No Chain 0.5767 0.5081 0.5424 0.5410 0.5310 0.5686
LLaVA-Video Chain 0.6292 0.5828 0.5875 0.5895 0.5606 0.5592
No Chain 0.6137 0.5767 0.5754 0.5868 0.5518 0.5760
LLaVA-Video-Only Chain 0.5653 0.5047 0.5081 0.5236 0.4791 0.5155
No Chain 0.5491 0.4960 0.4993 0.5121 0.4717 0.4899
GPT-4o Chain 0.7100 0.6292 0.4623 0.6332 0.6669 0.7133
No Chain 0.7207 0.6063 0.5747 0.5713 0.6252 0.6521
Gemini-1.5-Flash Chain 0.7443 0.7026 0.7073 0.7052 0.6393 0.6534
No Chain 0.7564 0.7106 0.7052 0.7046 0.6380 0.6346
Table 4: Model performance for semantic and structural similarity metrics across different numbers of few-shot samples $k$. These are the raw results (metrics visualized in Figure 3 were normalized, as discussed in the metrics section). * refers to model reasoning traces that were less coherent (Appendix C.2). The highest-performance at each value of $k$ is bolded.
Metric Model k = 0 k = 1 k = 2 k = 4 k = 8 k = 16
Similariy-Trace VideoChat2 0.4138 0.3954 0.4084 0.4079 0.0734 0.0597
Video-ChatGPT 0.4484 0.4484 0.4533 0.4354 0.0266 0.0760
LongVA 0.4533 0.4865 0.4929 0.4994 0.4994 0.4998
LLaVA-Video 0.4915 0.4193 0.4552 0.4178* 0.4629* 0.4731*
GPT-4o 0.4631 0.4332 0.3437 0.4363 0.4479 0.4648
Gemini-1.5-Flash 0.5157 0.5060 0.5021 0.4891 0.4887 0.4850
Similarity-Step VideoChat 0.3864 0.3804 0.3856 0.3749 0.0909 0.0760
Video-ChatGPT 0.4524 0.4822 0.4836 0.4700 0.0578 0.1059
LongVA 0.3898 0.4148 0.4185 0.4231 0.4297 0.4310
LLaVA-Video 0.4462 0.4090 0.4330 0.4109 0.4641 0.4856
GPT-4o 0.4016 0.3540 0.2943 0.3574 0.3682 0.3837
Gemini-1.5-Flash 0.4340 0.4378 0.4311 0.4234 0.4231 0.4210
Similarity-Num Steps VideoChat2 0.5168 0.4623 0.5983 0.4764 0.0000 0.0000
Video-ChatGPT 0.3267 0.2053 0.2144 0.1989 0.0000 0.0000
LongVA 0.4186 0.4529 0.4246 0.4179 0.4164 0.4181
LLaVA-Video 0.7719 0.6970 0.9205 0.7762* 0.4167* 0.3750*
GPT-4o 0.6226 0.4480 0.3496 0.4575 0.4440 0.4584
Gemini-1.5-Flash 1.0376 0.9144 0.7312 0.6984 0.6436 0.6525
EmotionMetric VideoChat2 0.0754 0.2328 0.2530 0.2322 0.0034 0.0007
Video-ChatGPT 0.2358 0.1713 0.1904 0.0702 0.0000 0.0000
LongVA 0.3345 0.1306 0.3244 0.3075 0.3106 0.3121
LLaVA-Video 0.3579 0.1077 0.1038 0.2168* 0.1667* 0.2500*
GPT-4o 0.6502 0.2005 0.6631 0.6725 0.6577 0.6286
Gemini-1.5-Flash 0.4959 0.1218 0.4152 0.3782 0.3557 0.3755
DifferenceSequenceMetric VideoChat 0.2673 0.2690 0.2838 0.2659 0.2141 0.2144
Video-ChatGPT 0.2508 0.2352 0.2325 0.2335 0.2181 0.2111
LongVA 0.4266 0.4165 0.4102 0.4202 0.4045 0.3750
LLaVA-Video 0.3783 0.4819 0.4845 0.4907* 0.2589* 0.1754*
GPT-4o 0.4582 0.3956 0.3671 0.4167 0.4458 0.4638
Gemini-1.5-Flash 0.4591 0.5301 0.5243 0.5312 0.5240 0.5295
Table 5: Model Performance for metrics related to step-level evidence from modalities and external knowledge across different numbers of few-shot samples $k$. These are the raw results (metrics visualized in Figure 3 were normalized, as discussed in the metrics section). * refers to reasoning traces that were less coherent (Appendix C.2). The highest-performance at each value of $k$ is bolded.
Metric Model k = 0 k = 1 k = 2 k = 4 k = 8 k = 16
All Modality Steps VideoChat2 1.2328 1.2436 1.3573 1.1817 0.6357* 0.6184*
Video-ChatGPT 0.8526 0.6801 0.6684 0.6638 0.6289* 0.6519*
LongVA 2.1406 1.7335 1.6669 1.6440 1.5659 1.4525
LLaVA-Video 1.6148 2.2242 2.0526 1.9231* 0.7500* 0.5625*
GPT-4o 2.3245 1.7155 1.4566 1.7536 1.8676 1.9750
Gemini-1.5-Flash 2.4056 2.3806 2.3529 2.3146 2.2811 2.2798
Visual Steps VideoChat2 0.6097 0.4105 0.5821 0.6326 0.0168* 0.0000*
Video-ChatGPT 0.2677 0.0919 0.0774 0.0755 0.0027* 0.0000*
LongVA 1.1299 0.9246 0.8863 0.8883 0.7782 0.7442
LLaVA-Video 1.0301 0.9367 0.9269 0.9580* 0.0000* 0.0000*
GPT-4o 1.5894 1.4196 1.0725 1.4180 1.4922 1.5598
Gemini-1.5-Flash 1.3276 1.2806 1.1860 1.1395 1.0431 1.0513
Verbal Steps VideoChat2 0.8338 0.8836 0.8358 0.8351 0.6316* 0.6184*
Video-ChatGPT 0.6533 0.5831 0.5774 0.5777 0.6030* 0.6519*
LongVA 0.8096 0.5942 0.6528 0.5989 0.6464 0.6988
LLaVA-Video 0.7643 0.7880 0.7487 0.8042* 0.9167* 0.6875*
GPT-4o 0.7249 0.2811 0.3984 0.3057 0.3150 0.3320
Gemini-1.5-Flash 0.9316 0.7728 0.7709 0.7695 0.7360 0.7196
Vocal Steps VideoChat2 0.1810 0.1635 0.2066 0.0740 0.0000* 0.0000*
Video-ChatGPT 0.0578 0.0202 0.0188 0.0117 0.0000* 0.0000*
LongVA 0.5162 0.2288 0.1844 0.1359 0.1365 0.1243
LLaVA-Video 0.2486 0.2375 0.1269 0.0490* 0.0000* 0.0000*
GPT-4o 0.4302 0.1270 0.0589 0.1184 0.1270 0.1413
Gemini-1.5-Flash 0.4357 0.2984 0.2551 0.2291 0.2497 0.2579
External Knowledge Steps VideoChat2 0.0229 0.3546 0.3230 0.2402 0.0054* 0.0000*
Video-ChatGPT 0.0130 0.0000 0.0000 0.0000 0.0232* 0.0000*
LongVA 0.3156 0.5565 0.5175 0.7174 0.6259 0.5044
LLaVA-Video 0.0792 1.4661 1.6128 1.5734* 0.0000* 0.0000*
GPT-4o 0.6199 0.9108 0.7554 1.0402 1.1568 1.2366
Gemini-1.5-Flash 0.9856 1.4901 1.6190 1.6881 1.6854 1.6847
Num Steps VideoChat2 4.9455 4.7840 5.8526 4.7638 3.1158 2.8876
Video-ChatGPT 3.4316 2.8955 2.8985 3.4564 2.8888 3.0443
LongVA 3.4233 2.3573 2.1245 1.9455 1.8689 2.0475
LLaVA-Video 2.5574 4.0755 4.7333 4.4336* 2.2500* 2.5625*
GPT-4o 4.0921 4.0757 3.5935 3.7699 3.7583 3.8039
Gemini-1.5-Flash 5.1060 4.2197 4.4624 3.9747 4.0445 3.6854